OpenStudy (anonymous):

The length of each side of a square is represented by 2x-1, represent the perimeter of the square solve for x, please explain how you got ur answer and tell me what the lenght is and what is the perimeter is...thanks

OpenStudy (anonymous):

4(2x-1)

OpenStudy (anonymous):

or 8x-4=perimeter

OpenStudy (anonymous):

again length of all sides and perimeter are the same things... square has 4 sides so multiply 2x-1 by 4

OpenStudy (anonymous):

to get what sider1993 said

OpenStudy (radar):

P=4(2x-1)=8x-4 P=8x-4 8x=P+4 x=(P+4)/8

OpenStudy (anonymous):

yea, but my teacher says otherwise (manln)

OpenStudy (anonymous):

thank you radar

OpenStudy (anonymous):

what does your teacher say perimeter and lengths mean?